A random variable, Z, which has mean 0 and standard deviation 1, find (1) p(z>1.2) (2) p(-2.0<z<2.0) (3) p(-1.2<z<1.0)

1
Expert's answer
2022-02-01T16:49:57-0500

(1) "P(Z>1.2)=1-P(Z<1.2)=0.1151."


(2) "P(-2.0<Z<2.0)=P(Z<2)-P(Z<-2.0)=0.9544."


(3) "P(-1.2<Z<1.0)=P(Z<1.0)-P(Z<-1.2)=0.7262."